# Real Bers Embedding on the Line: Fisher-Rao Linearization, Schwarzian Curvature, and Scattering Coordinates

### v1

@@ 4.5. Asymptotics of $S_{p}$ as $p\to\infty$
From Definition 4.19, $S_{p}(\varphi)=(S+\frac{3}{2p}(u^{\prime})^{2})\mathrm{e}^{u/p}$. Expanding $\mathrm{e}^{u/p}=1+u/p+u^{2}/(2p^{2})+O(p^{-3})$ and collecting terms yields the expansion. The $W^{k,1}$ remainder estimate uses Fa i . Bruno and Leibniz as in Proposition 2.23. ∎
@@ 4.6. Sturm-Liouville interpretation
4.7. Semiclassical interpretation
Definition 4.28 (Semiclassical parameter).
Identify $p=1/\hbar$ where $\hbar>0$ is the semiclassical parameter. The $L^{p}$-Schwarzian becomes the $\hbar$-Schwarzian
$S_{\hbar}(\varphi):=S_{1/\hbar}(\varphi)=\left(S(\varphi)+\frac{3\hbar}{2}\left(\frac{\varphi^{\prime\prime}}{\varphi^{\prime}}\right)^{2}\right)\varphi^{\prime\hbar}.$
Corollary 4.29 (WKB expansion of $S_{\hbar}$).
As $\hbar\to 0$ (equivalently $p\to\infty$), Proposition 4.24 yields the full asymptotic expansion
$S_{\hbar}(\varphi)=S(\varphi)+\hbar S^{(1)}(\varphi)+\hbar^{2}S^{(2)}(\varphi)+O_{W^{k,1}}(\hbar^{3}),$
where, with $u=\log(\varphi^{\prime})$,
$S^{(1)}(\varphi)$ $=u\,S(\varphi)+\frac{3}{2}(u^{\prime})^{2},$
$S^{(2)}(\varphi)$ $=\frac{u^{2}}{2}\,S(\varphi)+\frac{3u(u^{\prime})^{2}}{2}.$
Each coefficient $S^{(n)}$ is a universal polynomial in $u$ and its derivatives, involving derivatives up to order $n+2$.
Proposition 4.30 (Semiclassical Liouville correction).
Apply Theorem 4.26 to the semiclassical Schrödinger operator $\hbar^{2}\partial_{y}^{2}+V(y)$ by writing it as $\hbar^{2}(\partial_{y}^{2}-q(y))$ with $q(y)=V(y)/\hbar^{2}$. Then
$\mathcal{L}_{\varphi}\circ\bigl(-\hbar^{2}\partial_{y}^{2}+V(y)\bigr)\circ\mathcal{L}_{\varphi}^{-1}=-\hbar^{2}\partial_{x}^{2}+V(\varphi(x))\varphi^{\prime}(x)^{2}-\frac{\hbar^{2}}{2}\,S(\varphi)(x).$
In particular, the Schwarzian contributes the $\hbar^{2}$-order correction term dictated by Liouville functoriality.
Proof.
Set $q(y)=-V(y)/\hbar^{2}$ in Theorem 4.26 and multiply the resulting identity by $\hbar^{2}$. ∎
@@ 5. Central Extensions and Cocycle Geometry
The Schwarzian derivative is the fundamental obstruction to lifting the composition law on diffeomorphisms to an affine structure. This obstruction is measured by Lie algebra cohomology and central extensions. The results in the preceding sections show that $L^{p}$-based coordinates linearize the right-invariant Finsler geometry on diffeomorphism and density spaces, yielding explicit distances and geodesics. In contrast, the projective and cohomological features governed by the Schwarzian and by central extensions do not admit an equally clean interpolation. On the one hand, the Bott-Thurston extension is rigid in the sense that the $p$-root Jacobian cocycle produces no new real-valued Lie algebra cohomology classes beyond the Gelfand-Fuchs class, up to an explicit scalar. On the other hand, the $L^{p}$-Schwarzian fails to satisfy a two-term cocycle identity: the unavoidable cross-term in its composition law shows that the projective curvature cannot be encoded by a simple $L^{p}$ cocycle.
A complementary phenomenon - and the main genuinely new ingredient of this section - is that although the Bott cocycle is not a group coboundary on $\operatorname{Diff}^{-\infty}(\mathbb{R})$, it becomes an explicit groupoid coboundary once one allows dependence on a density through its score. This produces a canonical transgression functional whose coboundary is exactly the Bott phase. The resulting mechanism provides a clean categorical “absorption” of the projective multiplier into a central extension of Koopman transport.
@@ 5.2. The Virasoro-Bott extension
is well-defined on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ under the standing decay assumptions. By definition, $\psi^{\prime}-1$ is integrable together with all derivatives, hence $\log(\psi^{\prime})$ and $(\log\psi^{\prime})^{\prime}=\psi^{\prime\prime}/\psi^{\prime}$ are integrable; similarly $\log(\varphi^{\prime}\circ\psi)$ is bounded with integrable derivatives, so the product is integrable and the integral converges. Normalization is immediate: $B(\operatorname{id},\psi)=0$ and $B(\varphi,\operatorname{id})=0$ since $\log(\operatorname{id}^{\prime})=0$ and $d\log(\operatorname{id}^{\prime})=0$.
@@ 5.3. $L^{p}$ cocycles and rigidity
The quantity $\varphi^{\prime 1/p}$ is the Jacobian cocycle for the natural action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $p$th-root densities. Variants of this Jacobian cocycle, and Bott-Thurston type constructions for general $p$, appear in the literature on $L^{p}$ diffeomorphism groups and density geometry; see in particular [24] for extensions of the Virasoro algebra by modules of $\lambda$-densities. The point here is that, in the present decay-controlled setting, one can work in explicit $p$-root coordinates and obtain a clean rigidity statement, namely the induced real-valued Lie algebra $2$-cocycle is always the Gelfand-Fuchs cocycle, up to the explicit scalar factor $p^{-2}$. In particular, there is no new real-valued Lie algebra $2$-cocycle coming from the symmetric pairing $\int u^{\prime}v^{\prime}\,\mathrm{d}x$, which is not alternating and therefore does not define a Chevalley-Eilenberg $2$-cochain.

### v2

@@ 3.5. Asymptotics of $S_{p}$ as $p\to\infty$
From Definition 3.19, $S_{p}(\varphi)=(S+\frac{3}{2p}(u^{\prime})^{2})\mathrm{e}^{u/p}$. Expanding $\mathrm{e}^{u/p}=1+u/p+u^{2}/(2p^{2})+O(p^{-3})$ and collecting terms yields the expansion. The $W^{k,1}$ remainder estimate uses Faà di Bruno and Leibniz as in Proposition 2.16. ∎
@@ 4. Central Extensions and Cocycle Geometry
We now develop the cohomological content of the Schwarzian obstruction through Lie algebra cohomology and central extensions.
The main results of this section are the nontriviality of the Gelfand-Fuchs class (Proposition 4.2), its integration to the Virasoro-Bott group extension via the Bott cocycle (Proposition 4.5), and the rigidity of $L^{p}$ cocycles (Theorem 4.7).
@@ 4.2. The Virasoro-Bott extension
is well-defined on $\operatorname{Diff}^{-\infty}(\mathbb{R})$ under the standing decay assumptions: by definition $\psi^{\prime}-1$ is integrable together with all derivatives, hence $\log(\psi^{\prime})$ and $(\log\psi^{\prime})^{\prime}=\psi^{\prime\prime}/\psi^{\prime}$ are integrable; similarly $\log(\varphi^{\prime}\circ\psi)$ is bounded with integrable derivatives, so the product is integrable and the integral converges. Normalization is immediate: $B(\operatorname{id},\psi)=0$ and $B(\varphi,\operatorname{id})=0$ since $\log(\operatorname{id}^{\prime})=0$ and $d\log(\operatorname{id}^{\prime})=0$.
@@ 4.3. $L^{p}$ cocycles and rigidity
The quantity $\varphi^{\prime 1/p}$ is the Jacobian cocycle for the natural action of $\operatorname{Diff}^{-\infty}(\mathbb{R})$ on $p$th-root densities. Variants of this Jacobian cocycle, and Bott-Thurston type constructions for general $p$, appear in the literature on $L^{p}$ diffeomorphism groups and density geometry. The point here is that, in the present decay-controlled setting, one can work in explicit $p$-root coordinates and obtain a clean rigidity statement, namely the induced real-valued Lie algebra $2$-cocycle is always the Gelfand-Fuchs cocycle, up to the explicit scalar factor $p^{-2}$. In particular, there is no new real-valued Lie algebra $2$-cocycle coming from the symmetric pairing $\int u^{\prime}v^{\prime}\,\mathrm{d}x$, which is not alternating and therefore does not define a Chevalley-Eilenberg $2$-cochain.